OpenStudy (anonymous):

Find the probability of drawing a pink card, not replacing it, and then drawing another pink card. 2 blue 3 green 1 yellow 4 pink

OpenStudy (anonymous):

\[\frac{4}{10}\times \frac{3}{9}\]

OpenStudy (amistre64):

\begin{array}l\color{red}{\text{s}}\color{orange}{\text{a}}\color{#9c9a2e}{\text{t}}\color{green}{\text{s}}\color{blue}{\text{ }}\color{purple}{\text{g}}\color{purple}{\text{o}}\color{red}{\text{t}}\color{orange}{\text{ }}\color{#9c9a2e}{\text{i}}\color{green}{\text{t}}\color{blue}{\text{;}}\color{purple}{\text{ }}\color{purple}{\text{e}}\color{red}{\text{a}}\color{orange}{\text{s}}\color{#9c9a2e}{\text{i}}\color{green}{\text{e}}\color{blue}{\text{ }}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{a}}\color{orange}{\text{s}}\color{#9c9a2e}{\text{i}}\color{green}{\text{e}}\color{blue}{\text{ }}\color{purple}{\text{a}}\color{purple}{\text{s}}\color{red}{\text{ }}\color{orange}{\text{p}}\color{#9c9a2e}{\text{i}}\color{green}{\text{}}\end{array}